# KEN KANGDA WREN1

### v6

@@ Introduction
Some model-risks of this type may be hedged or diversified away. Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
@@ 1.1 The Usual-Case Asset and Information Structure
    Any incremental change (if at all) to bullet-payoff $Y$ is such that the cumulative firmed-up (’cum-div’) asset-worth $\{Z_{t}\}$, with $Z_{T}\equiv Y$, follows the dynamics below, based on a standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to binary outcomes $\{B=b\}$ (e.g. ’high’ or ’low’) labelled by $b\in\mathcal{B}:=\{+,-\}$,
@@
    where its $B$-sure drifts $\{r^{Z}_{B,t}\}$ and $B$-independent volatilities $\{\sigma^{Z}_{t}\}$ are both $\{\mathcal{F}^{Z}_{t}\}$-predictable. The underlying $B$-sure probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space/measure $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with process $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$. The above allows the sequential testing of Wiener processes of differing drifts [^2] Text-book sequential tests (e.g. Peskir and Shirayev (2006)) concern processes of fixed drift and volatility. Our setup allows time-variations achieved through an absolutely continuous clock-change, or equivalently, $\{\mathcal{F}^{Z}_{t}\}$-predictable drifts and volatilities. For more background see Appendix A..
@@
    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible, firmed-up value, $\{Z_{t}\}$, exist/matter. By considering data that are exclusive to model-risk (price-irrelevant once model-sure), including when tangible-value data, $\{Z_{t}\}$, are absent or trivial, we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate, while non-model, routine, risks remain, independently evolving or unchanging.
@@ A Properties of Inferential Hypothesis Testing
    Standard Continuous-Time Inference. Given the setup and Item-5 of Section 1.1, the log-LR process of time-homogeneous Wiener data processes has the form below (Peskir and Shirayev (2006)), in terms of some constant $\sigma^{l}>0$ characterising the signal-to-noise ratio of the given data,
@@
    with $\{w_{\tau}\}$ denoting a standard Wiener process. Any time-variation may be introduced via an absolutely continuous clock-change $t\mapsto\tau(t)$. The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is again Wiener, but with continuous time-varying parameters (e.g. Kallsen (2006)). It and its resulting inferential belief process under Bayes’ Rule (3) read: in terms of some standard Wiener process $\{w_{t}\}$,
@@
    Regular Inference. Its defining feature is the predictable resolution of risky outcomes. In standard situations, with i.i.d data, risk-resolution occurs only at infinity. Predictable finite-time resolution can be modelled via a clock-change that brings resolution forward to some $T_{\mathcal{B}}\leq\infty$. The resulting log-LR process diverges as $t\to T_{\mathcal{B}}$, but the associated beliefs and so asset-pricing remain finite and continuous at $T_{\mathcal{B}}$ under Bayes’ Rule (3).
@@
    Non-Resolution and Novikov’s Condition. For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely and hence a finite total variance for the associated log-LR process of form (A.2). The resolution of $B$-outcomes (Item-5, Section 1.1) thus equates to the associated LR process (the exponentiation of (A.2)), or equivalently, (A.3), violating Novikov’s Condition, and vice versa.
@@
    Adjacency. Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to\infty$, since:
@@
    where the RHS is convergent under equivalence; this is the case even when each of $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ is resolving (i.e. divergent in total variance and so total drift).
@@
    No two resolving log-LR processes, unless identical up to a constant difference, can be deterministically and time-homogenously related to one another.
@@
    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$, so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; we then have $g^{\prime}(0)=1$ as $\{\hat{l}^{+/-}_{t}\}$ and $\{{l}^{+/-}_{t}\}$ are resolving. If $B=$-, so $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$, then almost-sure divergence $\lim_{t\to\infty}\hat{l}^{+/-}_{t}=-\infty=\lim_{t\to\infty}{l}^{+/-}_{t}$ requires the same.
@@ B PROOF OF PROPOSITION 1
Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under the reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d\hat{\mathbf{W}}^{T}_{B,t}=\hat{r}_{B,\text{ }t}dt+d\mathbf{W}^{T}_{-,t}$, $\hat{{r}}_{B,\text{ }t}dt=-d\hat{{RP}}^{B}_{t}$. To conclude, the RNE measure $\{S^{(0|A)}_{t}\}$ must have the canonical form: $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.

### v7

@@ Introduction
Some model-risks of this type may be replicated/hedged or diversified. Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
@@ 1.1 The Usual-Case Asset and Information Structure
    Let $Y$ be revealed continuously by a public record $\{Z_{t}\}$ of realised increments (decrements) to some known base, set to $\log 1=0$, where each $Z_{t}$ is the sum of such firmed-up changes to date; by setup $Y\equiv Z_{T}$. Let $\{Z_{t}\}$ be a diffusion, based on some standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to outcome $\{B=b\}$, $b\in\mathcal{B}:=\{+,-\}$,
@@
    where its $B$-sure drifts $\{r^{Z}_{B,t}\}$ and $B$-independent volatilities $\{\sigma^{Z}_{t}\}$ are both $\{\mathcal{F}^{Z}_{t}\}$-predictable. The underlying $B$-sure space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$. The above allows the sequential testing of Wiener processes of differing drifts [^2] Text-book sequential tests (e.g. Peskir and Shirayev (2006)) concern processes of fixed drift and volatility. Our setup allows time-variations achieved through an absolutely continuous clock-change, or equivalently, $\{\mathcal{F}^{Z}_{t}\}$-predictable drifts and volatilities. For more background see Appendix A..
@@
    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible firmed-up value, $\{Z_{t}\}$, exist/matter. By considering data exclusively relevant to model-risk (irrelevant once model-sure), we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate asset valuation and hence traders’ cognitive and trading activities, while non-model, routine, risks remain, independently evolving or hardly changing.
@@ A Properties of Inferential Hypothesis Testing
    Standard Continuous-Time Inference. Given setup (Section 1.1), the log-LR process for testing time-homogeneous Wiener data of drift $\mathbf{1}_{+}(B)r^{Z}$ and volatility $\sigma^{Z}$ has the form (Peskir and Shirayev (2006)): in terms of signal-to-noise $\sigma^{l}=\frac{r^{Z}}{\sigma^{Z}}$,
@@
    with $\{w_{\tau}\}$ being some standard Wiener process. Time-dependence obtains via an absolutely continuous clock-change $t\mapsto\tau(t)$: image $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is Wiener, with continuous time-varying parameters (e.g. Kallsen (2006)). It and the resulting belief process by Bayes’ Rule (3) read: in terms of a standard Wiener process $\{w_{t}\}$ in $t$-time,
@@
    Non-Resolution and Novikov’s Condition. For inference based on equivalent Wiener-measure pair ${\mathbf{W}}_{+}\sim\mathbf{W}_{-}$, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d{\mathbf{W}}_{+}|_{t}(\cdot)}{d\mathbf{W}_{-}|_{t}(\cdot)}<\infty$ and so a finite total variance for the log-LR process (A.2) almost surely. Risk-resolution (Item-5, Section 1.1) thus equates to the LR process, the exponentiation of (A.2), or equivalently (A.3), violating Novikov’s Condition, and vice versa.
@@
    Regular Inference. It resolves $B$-outcomes predictably and so continuously. With i.i.d data risk-resolution occurs only at $T_{\mathcal{B}}=\infty$. Finite-time resolution can be modelled via a clock-change that brings resolution forward. Note that log-LR processes (A.2) diverge as $t\to T_{\mathcal{B}}\leq\infty$ while the associated beliefs (A.3) stay finite and continuous. For assets with horizon or maturity $T<T_{\mathcal{B}}$ or with $T_{\mathcal{B}}\leq T<\infty$ but with model-risk $B$ not $\mathcal{F}^{Z}_{T}$-measurable, our model-risk pricing applies all the same on $0,T_{\mathcal{B}}\wedge T)$ regardless. In the rest of the cases, where model-risk $B$ is $\mathcal{F}^{Z}_{T}$-measurable and hence replicable/hedgeable, its ’economic pricing’ (Remark 1) becomes irrelevant anyway.
@@
    Adjacency. Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a finite amount as $t\to\infty$, since:
@@
    where the RHS is bounded (Item-2), even if $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ are resolving (divergent).
@@
    No two resolving log-LR processes can be related to each other deterministically and time-homogenously without being identical up to a constant difference.
@@
    If $B=+$: $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; so $g^{\prime}(0)=1$ as $\lim_{t\to T_{\mathcal{B}}}\hat{l}^{+/-}_{t}=\infty=\lim_{t\to T_{\mathcal{B}}}{l}^{+/-}_{t}$. If $B=$-: $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$; so $g^{\prime}(0)=1$ as $\lim_{t\to T_{\mathcal{B}}}\hat{l}^{+/-}_{t}=-\infty=\lim_{t\to T_{\mathcal{B}}}{l}^{+/-}_{t}$.
@@ B PROOF OF PROPOSITION 1
Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under the reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d\hat{\mathbf{W}}^{T}_{B,t}=\hat{r}_{B,\text{ }t}dt+d\mathbf{W}^{T}_{-,t}$, $\hat{{r}}_{B,\text{ }t}dt=-d\hat{{RP}}^{B}_{t}$.